# thanks: Corresponding author

### v3

@@ I Introduction
In this work, we perform a systematic analysis of the masses of bound states in the toponium system, including the pseudoscalar state, $\eta_{t}$, and the vector state, $\psi_{t}$, as well as the mass of the triply top baryon, $\Omega_{ttt}$, within the framework of QCD sum rules. This approach, based on the QCD Lagrangian, provides a powerful tool for predicting hadronic properties, particularly in heavy quark systems, and its efficacy has been previously demonstrated in successful predictions of heavy hadron properties in agreement with experimental data Aliev:2009jt ; Aliev:2010uy ; Aliev:2012ru ; Agaev:2016mjb ; Azizi:2016dhy . Mass calculations within this framework require the use of a two-point correlation functions constructed from suitable interpolating currents, wherein the Operator Product Expansion (OPE), Borel transformation, and continuum subtraction based on quark–hadron duality are employed. In the present study, the OPE is considered up to dimension-eight non-perturbative condensates. The results of this analysis are expected to offer valuable theoretical insights for the interpretation of forthcoming LHC data and can inform future collider studies, such as those planned for CEPC and FCC-ee. The structure of the paper is as follows: Section II introduces the QCD sum-rule formalism and computational details; Section III discusses and compares the numerical results for the masses of $\eta_{t}$, $\psi_{t}$, and $\Omega_{ttt}$; finally, Section IV presents a summary and conclusions of this work. Alternatively: Relevant expressions are included in the Appendix.
@@ II.2 QCD side
In these relations, $\alpha$, $\beta$, $\delta$ denote Lorentz indices, and $f^{ABC}$ are the structure constants of the color group $\text{SU}_{c}(3)$, where $A,B,C=1,\,2\,\ldots 8$. In this analysis, the OPE is considered up to operators of dimension-eight. This expansion includes nonperturbative contributions associated with QCD vacuum condensates, namely $\langle G^{2}\rangle$ (dimension-four), $\langle G^{3}\rangle$ (dimension-six), and $\langle G^{2}\rangle^{2}$ (dimension-eight). Each term in the QCD side—containing both perturbative and nonperturbative contributions up to dimension-eight—has been derived following the comprehensive methodology presented in previous works Najjar:2024deh , where all relevant diagrams were calculated and analyzed. The final results are expressed as Fourier integrals and can be evaluated using standard mathematical techniques Najjar:2024deh ; Najjar:2024ngm . As mentioned earlier, on both the QCD and hadronic sides, the $\psi_{t}$ and $\Omega_{ttt}$ each possess two independent Lorentz structures. These structures are labelled as ($g_{\mu\nu}$, $q_{\mu}q_{\nu}$) and ($g_{\mu\nu}$, $\not\!qg_{\mu\nu}$), respectively. In the present work, the structure $g_{\mu\nu}$ is chosen for the vector toponium, and the corresponding structure $\not\!qg_{\mu\nu}$ is selected for $\Omega_{ttt}$. After applying the Borel transformation and subtracting the continuum, the final QCD-side expression of the correlation function for the studied particles is given as follows:
@@ III QCD Sum-Rule Mass Analysis
where $\mathrm{PC}$ denotes the pole contribution. The working intervals of the Borel parameter $M^{2}$ for the studied states are also summarized in Table 2. While the stability criteria mentioned above are satisfied for all channels, our numerical analysis reveals a characteristic structural pattern in the relative magnitudes of the OPE terms. In the baryonic channel $\Omega_{ttt}$, the nonperturbative contribution of dimension $D=4$ (the gluon condensate), amounting to approximately 64 %, dominates over the perturbative term ($D=0$), which contributes about 36 %. Similarly, in the mesonic (toponium) channels, a comparable anomaly arises at dimension $D=6$, whose contribution increases to roughly 61 %, surpassing the perturbative part. In light of this observed pattern, the convergence of the OPE remains valid—since the $D=8$ contributions are negligible—and the stability and consistency of the extracted results are ensured through the analysis of their dependence on $M^{2}.$
@@
respectively. This mass deficit reflects the presence of strong quantum correlations within the $t\bar{t}$ system. In the quantum-field picture, these correlations may be interpreted as possible manifestations of spin–spin entanglement between the top and anti-top quarks, which could play a role in enhancing the dynamical stability of the bound state. Such effects, if confirmed, would represent a distinctive quantum feature of the toponium system. In contrast to the two-body toponium states, the triply-top baryon ($\Omega_{ttt}$) is a three-quark system that must satisfy the color-singlet constraint, stabilized by strong inter-quark interactions. While the calculated central mass, $m_{\Omega_{ttt}}=517.81{}^{+1.82}_{-1.88}\,\mathrm{GeV}$, slightly exceeds the sum of the constituent top-quark masses, $\sum m_{t}=517.68{}^{+0.93}_{-0.93}\,\mathrm{GeV}$, the lower bound of the uncertainty range falls below this threshold. This behavior should not be interpreted as a physically negative binding energy, but rather as a reflection of the intrinsic uncertainties of the QCD sum-rule framework, including variations of the Borel mass and continuum threshold parameters. From a qualitative standpoint, one may speculate that quantum correlations or multi-body entanglement among the three top quarks could play a role in maintaining the internal colorless configuration and stabilizing the baryonic state, although a more detailed theoretical and experimental investigation is required to clarify these effects. The CMS collaboration has recently reported an excess of events near the $t\bar{t}$ threshold in final states with two charged leptons and multiple jets, using proton–proton collision data at $\sqrt{s}=13~\mathrm{TeV}$ corresponding to an integrated luminosity of 138 fb-1 CMS:2025kzt . This excess, observed with a statistical significance exceeding $5\sigma$, was interpreted as evidence for a quasi-bound pseudoscalar toponium state ($^{1}\mathrm{S}_{0}^{[1]}$}). The measured cross section relative to the fixed-order pQCD prediction is $8.8^{+1.2}_{-1.4}$ pb. No direct mass measurement of the pseudoscalar toponium was provided, though these results offer valuable guidance for theoretical studies. In parallel with the CMS results, the ATLAS Collaboration performed an independent and complementary analysis of $t\bar{t}$ production near the kinematic threshold using 140 fb-1 of proton–proton collision data at $\sqrt{s}=13$ TeV. The study focused on dileptonic final states with multiple jets and compared the observed invariant-mass spectrum to both a baseline perturbative QCD prediction and an extended model incorporating simulations of colour-singlet quasi-bound state formation. A statistically significant excess of events was observed over the baseline expectation, with a significance of 7.7 standard deviations, consistent with the production of an S-wave colour-singlet quasi-bound $t\bar{t}$ state as predicted by non-relativistic QCD. The corresponding excess cross section was measured to be $9.0\pm 1.3$ pb, further supporting the interpretation of near-threshold dynamics beyond pure perturbative descriptions. To facilitate a comprehensive comparison in light of recent $t\bar{t}$ measurements, Table 3 presents the mass estimates for $\eta_{t}$, $\psi_{t}$, and $\Omega_{ttt}$ from various theoretical studies, demonstrating notable consistency within the reported uncertainties. This overall agreement further reinforces the reliability of the present analysis.

### v4

@@ I Introduction
In this work, we perform a systematic analysis of the masses of bound states in the toponium system, including the pseudoscalar state, $\eta_{t}$, and the vector state, $\psi_{t}$, as well as the mass of the triply top baryon, $\Omega_{ttt}$, within the framework of QCD sum rules. This approach, based on the QCD Lagrangian, provides a powerful tool for predicting hadronic properties, particularly in heavy quark systems, and its efficacy has been previously demonstrated in successful predictions of heavy hadron properties in agreement with experimental data Aliev:2009jt ; Aliev:2010uy ; Aliev:2012ru ; Agaev:2016mjb ; Azizi:2016dhy . Mass calculations within this framework require the use of a two-point correlation functions constructed from suitable interpolating currents, wherein the Operator Product Expansion (OPE), Borel transformation, and continuum subtraction based on quark–hadron duality are employed. In the present study, the OPE is considered up to dimension-eight non-perturbative condensates. The results of this analysis are expected to offer valuable theoretical insights for the interpretation of forthcoming LHC data and can inform future collider studies, such as those planned for CEPC and FCC-$e$. The structure of the paper is as follows: Section II introduces the QCD sum-rule formalism and computational details; Section III discusses and compares the numerical results for the masses of $\eta_{t}$, $\psi_{t}$, and $\Omega_{ttt}$; finally, Section IV presents a summary and conclusions of this work. Alternatively: Relevant expressions are included in the Appendix.
@@ II.2 QCD side
In these relations, $\alpha$, $\beta$, $\delta$ denote Lorentz indices, and $f^{ABC}$ are the structure constants of the color group $\text{SU}_{c}(3)$, where $A,B,C=1,\,2\,\ldots 8$. In this analysis, the OPE is considered up to operators of dimension-eight. This expansion includes nonperturbative contributions associated with QCD vacuum condensates, namely $\langle G^{2}\rangle$ (dimension-four), $\langle G^{3}\rangle$ (dimension-six), and $\langle G^{2}\rangle^{2}$ (dimension-eight). Each term in the QCD side—containing both perturbative and nonperturbative contributions up to dimension eight—has been explicitly evaluated in the present work. For specific technical details related to the expansion in terms of Feynman diagrams and standard QCD sum-rule procedures, we refer the reader to Ref Najjar:2024deh . The final results are expressed as Fourier integrals and can be evaluated using standard mathematical techniques Najjar:2024deh ; Najjar:2024ngm . As mentioned earlier, on both the QCD and hadronic sides, the $\psi_{t}$ and $\Omega_{ttt}$ each possess two independent Lorentz structures. These structures are labelled as ($g_{\mu\nu}$, $q_{\mu}q_{\nu}$) and ($g_{\mu\nu}$, $\not\!qg_{\mu\nu}$), respectively. In the present work, the structure $g_{\mu\nu}$ is chosen for the vector toponium, and the corresponding structure $\not\!qg_{\mu\nu}$ is selected for $\Omega_{ttt}$. After applying the Borel transformation and subtracting the continuum, the final QCD-side expression of the correlation function for the studied particles is given as follows:
@@ III QCD Sum-Rule Mass Analysis
where $\mathrm{PC}$ denotes the pole contribution. The working intervals of the Borel parameter $M^{2}$ for the studied states are also summarized in Table 2. While the stability criteria mentioned above are satisfied for all channels, our numerical analysis reveals a characteristic structural pattern in the relative magnitudes of the OPE terms. In heavy-quark systems, the large top-quark mass establishes the dominant physical scale. This significantly enhances the correlator’s sensitivity to non-perturbative vacuum fields. The correlation function, expressed via propagators, generates non-perturbative terms across various mass dimensions by Eq. (II.2). Within the working Borel window, the substantial top-quark mass, which appears explicitly in the correlation function, enhances the gluon condensate contributions. Consequently, these contributions achieve numerical significance without compromising the convergence of the OPE. In the baryonic channel $\Omega_{ttt}$, the nonperturbative contribution of dimension $D=4$ (the gluon condensate), amounting to $63.7\%$, dominates over the perturbative term ($D=0$), which contributes $35.8\%$. Similarly, in the mesonic (toponium) channels, a comparable anomaly arises at dimension $D=6$, whose contribution increases to $61.2\%$, surpassing the perturbative part. In light of this observed pattern, the convergence of the OPE remains valid—since the dimension $D=8$ contributions are negligible (about $0.09\%$ in both cases)—and the stability and consistency of the extracted results are ensured through the analysis of their dependence on $M^{2}$. To demonstrate the convergence of the OPE terms, Fig. 1 has been generated for triply- top baryon as an example. The left panel of Fig. 1 displays the relative contributions of successive OPE terms with dimensions $D=4,6,8$ as a function of the Borel parameter (at the central value of $s_{0}$). The right panel provides a magnified view of the contributions from terms with $D=6,8$ , highlighting the specific impact of each of these operators on the convergence.
Figure 1: Left: Dependence of OPE terms contributions ($D=4,6,8$) for the triply-top baryon ($\Omega_{ttt}$) on the Borel mass parameter ($M^{2}$) at thecentral value of the continuum threshold ($s_{0}$). Right : Dependence of OPE terms contributions ($D=6,8$) for the triply-top baryon ($\Omega_{ttt}$) on the Borel mass parameter ($M^{2}$) at thecentral value of the continuum threshold ($s_{0}$).
@@
respectively. This mass deficit may reflect the presence of strong quantum correlations within the $t\bar{t}$ system. In contrast to the two-body toponium states, the triply-top baryon ($\Omega_{ttt}$) is a three-quark system that must satisfy the color-singlet constraint, stabilized by strong inter-quark interactions. While the calculated central mass, $m_{\Omega_{ttt}}=517.81{}^{+1.82}_{-1.88}\,\mathrm{GeV}$, slightly exceeds the sum of the constituent top-quark masses, $\sum m_{t}=517.68{}^{+0.93}_{-0.93}\,\mathrm{GeV}$, the lower bound of the uncertainty range falls below this threshold. This behavior should not be interpreted as a physically negative binding energy, but rather as a reflection of the intrinsic uncertainties of the QCD sum-rule framework, including variations of the Borel mass and continuum threshold parameters. From a qualitative standpoint, one may speculate that quantum correlations among the three top quarks could play a role in maintaining the internal colorless configuration and stabilizing the baryonic state, although a more detailed theoretical and experimental investigation is required to clarify these effects. . The CMS collaboration has recently reported an excess of events near the $t\bar{t}$ threshold in final states with two charged leptons and multiple jets, using proton–proton collision data at $\sqrt{s}=13~\mathrm{TeV}$ corresponding to an integrated luminosity of 138 fb-1 CMS:2025kzt . This excess, observed with a statistical significance exceeding $5\sigma$, was interpreted as evidence for a quasi-bound pseudoscalar toponium state ($^{1}\mathrm{S}_{0}^{[1]}$}). The measured cross section relative to the fixed-order pQCD prediction is $8.8^{+1.2}_{-1.4}$ pb. No direct mass measurement of the pseudoscalar toponium was provided, though these results offer valuable guidance for theoretical studies. In parallel with the CMS results, the ATLAS Collaboration performed an independent and complementary analysis of $t\bar{t}$ production near the kinematic threshold using 140 fb-1 of proton–proton collision data at $\sqrt{s}=13$ TeV. The study focused on dileptonic final states with multiple jets and compared the observed invariant-mass spectrum to both a baseline perturbative QCD prediction and an extended model incorporating simulations of colour-singlet quasi-bound state formation. A statistically significant excess of events was observed over the baseline expectation, with a significance of 7.7 standard deviations, consistent with the production of an S-wave colour-singlet quasi-bound $t\bar{t}$ state as predicted by non-relativistic QCD. The corresponding excess cross section was measured to be $9.0\pm 1.3$ pb, further supporting the interpretation of near-threshold dynamics beyond pure perturbative descriptions. To facilitate a comprehensive comparison in light of recent $t\bar{t}$ measurements, Table 3 presents the mass estimates for $\eta_{t}$, $\psi_{t}$, and $\Omega_{ttt}$ from various theoretical studies, demonstrating notable consistency within the reported uncertainties. This overall agreement further reinforces the reliability of the present analysis.